OpenStudy (anonymous):

A chemist mixes some 70% solution with some 40% solution to obtain 120 gallons of 50% solution. Find the number of gallons of each solution. A. define variables for unknown B. write a system of equations to model the problem C. Solve System D. Give answer in a sentence

OpenStudy (dumbcow):

 ---> x + y = 120 ---> 70x + 40y = 50(120)

OpenStudy (anonymous):

A. unknown variables : x + y = 120 ? B. system of equations : 70x + 40y = 50 (120) ? @dumbcow

OpenStudy (dumbcow):

do you know what a variable is?

OpenStudy (anonymous):

x and y

OpenStudy (dumbcow):

very good... those are variables because they represent unknown values, namely the number of gallons for 70%solution and 40%solution a system of equations implies more than 1 equation we have 2 equations in our system

OpenStudy (anonymous):

and the equations are x + y = 120 70x + 40y = 50 (120)

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

solve by using substitution x+y = 120 ---> y = 120 - x 70x + 40(120-x) = 50(120) solve single equation for x

OpenStudy (anonymous):

x = 70?

OpenStudy (dumbcow):

nope

OpenStudy (anonymous):

1.25?

OpenStudy (anonymous):

I did 70 (0) + 40 (120-0) = 50 ( 120)

OpenStudy (dumbcow):

?? that is assuming x=0 , not finding what x actually is

OpenStudy (anonymous):

I thought thats what substituting was

OpenStudy (anonymous):

how do i find x then?

OpenStudy (dumbcow):

i kinda helped by doing the substitution for you i solved for y in 1st equation ----> y = 120-x and plugged it into 2nd equation ..... replaced "y" with "120-x" think back to when you first learned algebra, you want to isolate the variable step 1: distribute the 40(120-x) step 2: add like terms step 3: get "x" term on left side, constant on right side step 4: divide

OpenStudy (anonymous):

oh yea that makes sense now